OpenStudy (anonymous):

PLEASE HELP IM GONNA FAIL MATH Explain the procedure when using the elimination method to get rid of one of the variables

OpenStudy (firejay5):

So explain how to do elimination

OpenStudy (anonymous):

I'll give u a medal if u just pls give methe anwser

OpenStudy (anonymous):

you must try to get one of the variables to equal 0

OpenStudy (anonymous):

do you need an example?

OpenStudy (firejay5):

For example if you had 4a + 2b = 15 2a + 2b = 7 you would need to get the B's to equal zero in this case

OpenStudy (anonymous):

if your dealing with systems of equations: http://www.wyzant.com/Help/Math/Precalculus/Systems_Of_Equations/Elimination_Method.aspx

OpenStudy (firejay5):

so you would (2a + 2b = 7) *-1 and it would be this: -2a - 2b = -7

OpenStudy (firejay5):

the B's now cancel out. 4a - -2a = 6a = 8 <---- 15 - 7 = 8

OpenStudy (firejay5):

the answer would actually be 2a 4a + -2a = 2a = 22<---- 15 + 7 = 22. 2a = 22 then divide 2 from both sides to get A = 11. Then insert 11 into both equations and solve to get y